OpenStudy (idealist10):

Find a second solution y2 for (x^2-4)y"+4xy'+2y=0; y1=1/(x-2) that isn't a constant multiple of the solution y1.

OpenStudy (idealist10):

@SithsAndGiggles

OpenStudy (anonymous):

Set \(y_2=\dfrac{u}{x-2}\), then \[{y_2}'=\frac{u'}{x-2}-\frac{u}{(x-2)^2}\quad\text{and}\quad{y_2}''=\frac{u''}{x-2}-\frac{2u'}{(x-2)^2}+\frac{2u}{(x-2)^3}\] So you have the ODE \[\begin{align*} 0&=(x^2-4)y''+4xy'+2y\\\\ 0&=(x^2-4)\left(\frac{u''}{x-2}-\frac{2u'}{(x-2)^2}+\frac{2u}{(x-2)^3}\right)+4x\left(\frac{u'}{x-2}-\frac{u}{(x-2)^2}\right)\\ &\quad\quad+2\frac{u}{x-2}\\\\ 0&=(x+2)u''-\frac{2(x+2)}{x-2}u'+\frac{2(x+2)}{(x-2)^2}u+\frac{4x}{x-2}u'-\frac{4x}{(x-2)^2}u+\frac{2}{x-2}u\\\\ 0&=(x+2)u''+\frac{2x-4}{x-2}u'\end{align*}\]

OpenStudy (anonymous):

Missed some simplification, the ODE is equivalent to \[(x+2)u''+2u'=0\] From here you can replace \(u'=z\) and you have an ODE linear in \(z\): \[(x+2)z'+2z=0\]

OpenStudy (idealist10):

I got the right answer! Thanks!